A Toboggan loaded with students (total weight 150N) slides down a snow covered slope. The at constant angle 45, and the toboggan is so well waxed that there is hill slopes a virtually no friction, What is its acceleration?
Answer #1

From the given conditions

F parallel = m*g*sin 45

F perpendicular= m*g* cos 45

There is no other force parallel to the incline to counteract the parallel component of gravity .so net force is equal to the parallel force.

m*a = m*g*sin 45

a = g*sin45

a= 9.8*sin45

a= 6.92 m/s^2